OpenStudy (anonymous):

Solve sec^2(x)=1 over the interval [(-pi/2), (5pi/2)]

OpenStudy (amistre64):

sec = 1/cos so when does cos(x) = 1 or -1 in the interval?

OpenStudy (anonymous):

is y supposed to equal to 1 or x?

OpenStudy (anonymous):

\[\sec^2(x)=1\] \[\sec(x)=1\] or \[\sec(x)=-1\] \[\frac{1}{\cos(x)}=1\] or \[\frac{1}{\cos(x)}=-1\]

OpenStudy (anonymous):

which is the same thing as saying \[\cos(x)=1\] so \[x=0,x=2\pi\] or \[\cos(x)=-1\] \[x=\pi\]